Show that T:R^3 →R^2: T(x,y,z)= (2x+y-z,x+z) is a linear transformation. Verify that T satisfies the Rank-nullity theorem


1
Expert's answer
2021-07-05T14:40:25-0400

Let "\alpha=(x_1,y_1,z_1)\isin R^3,

\beta=(x_2,y_2,z_2)\isin R^3"

Then "T(\alpha)=(2x_1+y_1-z_1,x_1+z_1)"

"T(\beta)=(2x_2+y_2-z_2,x_2+z_2)"

"\alpha+\beta=(x_1+x_2,y_1+y_2,z_1+z_2)"

"T(\alpha+\beta)=(2(x_1+x_2)+(y_1+y_2)-(z_1+z_2),(x_1+x_2)+(z_1+z_2))"

"=((2x_1+y_1-z_1)+(2x_2+y_2-z_2),(x_1+z_1)+(x_2+z_2))"

"=T(\alpha)+T(\beta)"

Also let "c\isin R." 

Then

 "c\alpha=(cx_1,cy_1,cz_1)"

"T(c\alpha)=(2cx_1+cy_1-cz_1,cx_1+cz_1)"

"=c(2x_1+y_1-z_1,x_1+z_1)"

"=cT(\alpha)"

Thus 

"T(\alpha+\beta)=T(\alpha)+T(\beta)" 

for all "\alpha,\beta\isin R^3"

and 

"T(c\alpha)=cT(\alpha)" for all "c\isin R" and "\alpha\isin R^3."

Hence "T" is a linear transformation.


"K""er" "T=" {"(x,y,z)\isin R^3:T(x,y,z)=(0,0)"}

Let "(x_1,y_1,z_1)\isin" "Ker" "T"

Then "2x_1+y_1-z_1=0,x_1+z_1=0."

From second equation we get "x_1=-z_1"

Let "x_1=k" (say) , then "z_1=-k"

Now from first equation we get , "y_1=-3k"

Therefore "(x_1,y_1,z_1)=k(1,-3,-1)=c(-1,3,1)""c\isin R"

Let "\alpha=(-1,3,1)". Then "Ker" "T="linear span of "\alpha=L"{"\alpha" } and "dim""Ker" "T=1."


"ImT" is the linear span of the vectors "T(\alpha_1),T(\alpha_2),T(\alpha_3)" where {"\alpha_1,\alpha_2,\alpha_3"} is any basis of "R^3".

{"\epsilon_1=(1,0,0),\epsilon_2=(0,1,0),\epsilon_3=(0,0,1)"} is a basis of "R^3".

Now "T(\epsilon_1)=(2,1),T(\epsilon_2)=(1,0),T(\epsilon_3)=(-1,1)."

"ImT=L"{"(2,1),(1,0),(-1,1)" }

These vectors are linearly dependent in "R^2."

But the subset {"(2,1),(1,0)"} is linearly independent in "R^2".

"\therefore ImT=L"{"(2,1),(1,0)" }

Therefore "dim" "ImT=2."


Now according to the Rank- Nullity theorem we know that

"dim" "KerT+" "dim" "ImT" "=dim" "R^3"

Here, "dim" "KerT+dim" "ImT=1+2=3".

"dim" "R^3=3"

Therefore "T" satisfies the Rank-Nullity theorem.